OpenStudy (anonymous):

Solve: 7(x − 1) = 5(x + 2) + 1 9 18 2 4

OpenStudy (***[isuru]***):

hi, 7(x − 1) = 5(x + 2) + 1 7x - 7 = 5x + 10 +1 7x -5x = 10 + 1 +7 2x = 18 x = 9 hope this will help ya out!!

OpenStudy (***[isuru]***):

did ya get it ?